OpenStudy (anonymous):

Circle O is defined by the equation x2 + (y - 2)2 = 25. Plot the center of Circle O and type in the coordinates of one point with integral values that lies on Circle O.

OpenStudy (anonymous):

so th radius would be 5

OpenStudy (ikram002p):

the circle equation of center(x_1,y_1) and raduis r \(\large (x-x_1)^2+(y-y_1)^2=r^2\)

OpenStudy (ikram002p):

yes correct :) what about center?

OpenStudy (anonymous):

I know the y would 4 right?? and x i dont know.

OpenStudy (ikram002p):

u can write the equation like this no y \(\neq \) 4 , check again \(\large (x-0)^2+(y-2)^2=r^2\)

OpenStudy (anonymous):

-4??

OpenStudy (ikram002p):

\((x-0)^2+(y-2)^2=5^2\) how did u get 4 ?

OpenStudy (anonymous):

-2 ^2 equals 4

OpenStudy (ikram002p):

but it should not be -2^2 it only be 2 , as the example before

OpenStudy (anonymous):

oh i see,

OpenStudy (ikram002p):

so what is the center and raduis again ?

OpenStudy (anonymous):

0,2 and the radius 5

OpenStudy (ikram002p):

correct :) and draw the circle

OpenStudy (anonymous):



OpenStudy (ikram002p):



OpenStudy (anonymous):

okay! thanks that was helpful, really helpful!!

OpenStudy (anonymous):

i don't quite get this!

OpenStudy (anonymous):

hahah sorry :(

OpenStudy (ikram002p):

but we dint finish the first one yet :o

OpenStudy (anonymous):

oh, i thought we did xD continue sorry.

OpenStudy (ikram002p):

lol sorry nvmits done :P

OpenStudy (ikram002p):

ok the second one m<SOT= 2 * m<SRT

OpenStudy (anonymous):

yeah,

OpenStudy (ikram002p):

find , m <SRT :o m <SRT=m<SOT / 2=68/2

OpenStudy (anonymous):

sorry im so out of it, i have slept a total of 5 hrs in the last 3days and i'm trying really hard to get these answers so i can rest.

OpenStudy (ikram002p):

how old are u ?

OpenStudy (anonymous):

15

OpenStudy (ikram002p):

ohh sleep more xD nw for the question m <SRT=m<SOT / 2=68/2 so m<SRT=68/2=34

OpenStudy (anonymous):

srt = 34?

OpenStudy (ikram002p):

yep